OpenStudy (anonymous):

find the mean, median, and mode of the data set. 1. 12 13 14 9 7 8 15 10 11 6

OpenStudy (anonymous):

@ash2326 @campbell_st

OpenStudy (anonymous):

First put them in order.

OpenStudy (anonymous):

Then to find the median find the middle number, in this case there will be two...so add them together and divide by 2. This will be your median Mode: the number that appears the most often Mean add up all the digits and divide by 10. This is your mean. Hope i helped though I didn't give the answer:)

OpenStudy (campbell_st):

median... arrange them in order... smallest to largest... the median is a the middle. the median has exactly half way between the 5th and 6th terms mode... most frequent score mean.... total the scores then divide by the number of scores

OpenStudy (cwrw238):

arranging in ascending order 6 7 8 9 10 11 12 13 14 15 the middle 2 are 10 and 11 so median = (10 + 11) / 2 = 21/2 = 10.5

OpenStudy (anonymous):

whats the mean and mode?

OpenStudy (cwrw238):

no mode add all the numbers and divide by 10 to get the mean

OpenStudy (cwrw238):

all numbers are different so theres no mode

OpenStudy (anonymous):

@cwrw238 10.5,10.5,none?

OpenStudy (cwrw238):

yes